LBC Corporation makes and sells a product called Product WZ. Each unit of Product WZ requires 3.0 hours of direct labor at the rate of $26.00 per direct labor-hour. Management would like you to prepare a Direct Labor Budget for June. The company plans to sell 49,000 units of Product WZ in June. The finished goods inventories on June 1 and June 30 are budgeted to be 610 and 170 units, respectively. Budgeted direct labor costs for June would be:

Multiple Choice:

  • $3,825,180

  • $3,787,680

  • $3,806,430

  • $1,266,500
Answer #1

Number of units to be produced = Budgeted sales + Ending inventory - Beginning inventory

= 49,000 + 170 - 610

= 48,560 units

Time required for each unit = 3 hours

Rate per direct labour hour = $26

Direct labour cost per unit = 3*26 = $78

Budgeted direct labour cost for june = 48,560 units * $78 = $37,87,680

Therefore, Budgeted labour cost for june = $37,87,680